### v1

# LLMs in the Heart of Differential Testing: A Case Study on a Medical Rule Engine

Abstract.
The Cancer Registry of Norway (CRN) uses an automated cancer registration support system (CaReSS) to support core cancer registry activities, i.e, data capture, data curation, and producing data products and statistics for various stakeholders. GURI is a core component of CaReSS, which is responsible for validating incoming data with medical rules. Such medical rules are manually implemented by medical experts based on medical standards, regulations, and research. Since large language models (LLMs) have been trained on a large amount of public information, including these documents, they can be employed to generate tests for GURI. Thus, we propose an LLM-based test generation and differential testing approach (LLMeDiff) to test GURI. We experimented with four different LLMs, two medical rule engine implementations, and real medical rules to investigate the hallucination, success, time efficiency, and robustness of the LLMs to generate tests, and these tests’ ability to find potential issues in GURI. Our results showed that GPT-3.5 hallucinates the least, is the most successful, and is generally the most robust; however, it has the worst time efficiency. Our differential testing revealed 22 medical rules where implementation inconsistencies were discovered (e.g., regarding handling rule versions). Finally, we provide insights for practitioners and researchers based on the results.
1. Introduction
The cancer registration support system (CaReSS) is a real-world socio-technical software system developed and operated by the Cancer Registry of Norway (CRN), a public organization in Norway under the National Institute of Public Health (NIPH). CaReSS receives cancer patient data through cancer messages from other public and private entities, including hospitals, laboratories, and health registries (e.g., national patient registry). Such data is then processed by CaReSS to produce outputs (including data and statistics) for its end users, such as the public, researchers, and policymakers.
One core functionality of CaReSS is to validate incoming cancer messages and medical data with a dedicated rule validation software system called GURI, which checks the correctness of received cancer messages, validates the data with various medical rules, aggregates individual cancer messages into cancer cases, and validates the aggregation. The rules are manually defined by medical coders based on various standards and regulations, such as International Classification of Diseases (ICD)-10 (10th edition), ICD-O-2 (ICD for Oncology, 2nd edition) for solid tumors, and ICD-O-3 (ICD for Oncology, 3rd edition) for non-solid tumors. Extensively testing GURI is important since an incorrect implementation would produce erroneous data or statistcs, which are used in medical research, monitoring and evaluation of cancer care.
As large language models (LLMs) have been trained on a large amount of public information, including the standards, regulations, and other documents that are the sources for specifying the medical rules, we employ these LLMs to generate tests for these medical rules. Consequently, we developed an LLM-based differential testing approach called LLMeDiff to find potential faults in GURI. An LLM is used to generate tests from medical rules (that yield Pass, Fail, and NotApplied results for each rule), which are then executed on GURI and a simple reference implementation (Dvare++). A mismatch between the outputs from these two flags a potential inconsistency stemming from an implementation issue in GURI.
In our previous works, we studied a state-of-the-art test generation tool (i.e., EvoMaster [3]) applied to GURI [26], reduced its number of executed tests with machine learning (ML) classifiers [21], and built cyber-cyber digital twins of GURI [29]. In contrast, in this paper, we improve on our previous works by (1) leveraging LLMs to generate valid medical data as test inputs, (2) addressing the oracle problem of rule results through differential testing with a reference implementation, and (3) reducing the number of executed tests against GURI by only generating three tests per rule.
We experimented LLMeDiff by selecting four LLMs (i.e., Mistral, Llama2, Mixtral, and GPT-3.5), real medical rules and their mutations, and two rule engine implementations (i.e., GURI and Dvare++). We focused on two perspectives: (1) how effective, efficient, and robust the selected LLMs are in generating tests from medical rules; and (2) how many potential issues we can identify in the implementation of GURI with our differential testing approach.
Our results showed that LLMs are highly effective in generating medical rule tests. In particular, GPT-3.5 is the most effective LLM with hardly any hallucinations (i.e., mean completion rate of %) and being successful in generating Pass (%), Fail (%), and NotApplied (%) tests. However, GPT-3.5 is also the least efficient model with compared to the fastest models, i.e., Mixtral with and Mistral with . In terms of robustness, GPT-3.5 performs the best for Pass tests with % and equally to Mixtral and Mistral for NotApplied tests with %. Only for Fail tests, GPT-3.5 with % is inferior to Mistral and Mixtral with both having %. Finally, regarding the differential testing results, we observe that Mixtral has the highest number of matches, while Mistral has the highest number of mismatches per rule. A total of 44 rule mismatches between GURI results and Dvare++ for Mistral, followed by GPT-3.5 with 32, Llama2 with 29 and Mixtral with 27. While not all mismatches can be considered faults, our analysis reveals 22 rules that show inconsistencies in terms of handling either rule versions, date format, or variable dependencies.
These results suggest that LLMs are effective rule test generators to discover mismatches with differential testing, improving the current state of practice at the CRN. Whether a mismatch is an actual fault still requires manual root cause analysis by the CaReSS developers.
2. Real-World Application Context
The CRN compiles information on cancer patients, covering data related to diagnosis and medical treatments sourced from various healthcare organizations (e.g., hospitals and laboratories) as cancer messages. The CRN created CaReSS to analyze and validate cancer messages and offer decision support to stakeholders like policy makers and researchers. The accuracy of the decisions made with CaReSS relies on the validity of the cancer messages it stores. To validate cancer messages, the CRN developed a subsystem of CaReSS named GURI—our application context. GURI is a web-based system that provides several representational state transfer (REST) application programming interface (API) endpoints corresponding to various functionalities. One of the endpoints is the validation endpoint (/api/messages/validation) for validating cancer messages against medical rules.
Dvare111https://github.com/dvare/dvare-framework/ is a rule-expression language engine that runs the rules on the data and asserts the results into true or false depending on if any of the condition matches. GURI employs Dvare as the underlying engine to evaluate cancer messages against medical rules. While Dvare yields a Boolean result, GURI categorizes the results based on logical operations, particularly the implies operator. Consider an example rule with two operands joined by an implies operation, as illustrated in Fig. 1. In this scenario, GURI handles the rules as follows: (1) If the LeftOperand is true, the rule is applied and can be either Pass or Fail depending on the RightOperand (true for Pass, and false for Fail). (2) If the LeftOperand is false, the rule is NotApplied, irrespective of the RightOperand. This categorization enables GURI to present rule validation outcomes with enhanced clarity for analysis and prediction within CaReSS.
Testing GURI is essential to ensure the accuracy of the cancer data and statistics produced by CaReSS [21]. The challenges of testing GURI are as follows. (1) Evolution: GURI and Dvare continuously evolve in response to changing cancer messages format and medical rules dictated by standards/regulations or through cancer research. (2) High-cost: Testing multiple evolving implementations of GURI with various underlying Dvare versions incurs high time and resource costs. (3) Labor-intensive: Manual testing of each version GURI in different development phases is a laborious task. Therefore, a cost-effective and automated approach is necessary to test the evolving GURI [27]. (4) Absence of Test Oracles: For medical rule testing, there are no precise test oracles to determine whether the rules are implemented correctly, posing additional challenges for automated testing.
3. Approach
Figure 2 shows LLMeDiff’s overview in two stages: (1) medical rule test generation and (2) differential testing of GURI.
Test Generation
This stage uses an LLM, which takes a prompt and a set of rules as input. LLMeDiff executes one prompt for each rule, resulting in three tests per rule: one Pass, one Fail, and one NotApplied test. A test has (only) the variable values that let the rule Pass, Fail, or NotApplied.
LABEL:lst:prompt shows the prompt divided into a system and user part.
In the system part (system_msg), we set up the LLM to follow these steps:
fnum@@desciitemStep 1: be precise;
fnum@@desciitemStep 2: identify rule variables and generate tests:
(1) Pass(satisfying_case),
(2) Fail(violating_case),
and
(3) NotApplied(invalid_case);
fnum@@desciitemStep 3: do not be repetitive;
fnum@@desciitemStep 4: express your confidence (confidence_score);
and
fnum@@desciitemStep 5: generate a JavaScript Object Notation (JSON) object as the output.
Note that we use different terms in the prompt, as the rules are similar to constraints in propositional logic and, hence, we expect the LLM to understand these terms better than the terms specific to GURI.
In the user part (user_msg), we instruct the LLM to generate the three tests for a given rule, for which the prompt can be parameterized ({rule}).
Differential Testing
In this stage, we execute the tests against two system under tests (SUTs):
{inparablank}
4. Experimental Evaluation
To evaluate LLMeDiff, we perform a laboratory experiment [37] on four LLMs and real-world medical rules from the CRN.
4.1. Research Questions
Our experiment investigates the following research questions (RQs).
RQ1:
How effective and efficient are LLMs in generating tests from medical rules?
RQ2:
How robust are LLMs when generating tests?
RQ3:
To what extent can the generated tests be used for differential testing of a real-world medical rule engine?
RQ1 studies the differences in the abilities of the selected LLMs in generating tests from three aspects:
(1) hallucinations, i.e., syntactically invalid outputs;
(2) successful generations, i.e., semantically correct outputs (e.g., a test that should pass actually passes);
and
(3) time efficiency.
RQ2 focuses on the robustness of the selected LLMs to study their trustworthiness when generating tests for slightly mutated rules.
RQ3 is designed to investigate the effectiveness of the generated tests in evaluating the GURI system, thereby studying the applicability of LLMeDiff to a real-world system.
4.2. LLM Selection and Settings
We selected four LLMs:
(1) Mistral 7B denoted as Mistral [23],
(2) Llama 2 13B denoted as Llama2 [39],
(3) Mixtral 8x7b denoted as Mixtral [24],
and
(4) GPT-3.5 denoted as GPT-3.5 [33].
We selected these models while considering availability, diversity, performance across general-purpose tasks, and popularity among the AI community.
GPT-3.5 from OpenAI, is a generative pre-trained transformer, while Llama2, from Meta, is a pre-trained and fine-tuned LLM with 13 billion parameters. We choose Mistral from Mistral AI for its good performance in reasoning, mathematics, and code generation, outperforming Llama2 on multiple benchmarks [23]. Furthermore, Mixtral, a sparse mixture of experts model with open weights, distinguishes itself by matching or surpassing GPT-3.5 on various tasks [24]. Its impressive performance and widespread popularity within the AI community, particularly on platforms like HuggingFace222https://huggingface.co/ were the reasons to select it.
We used the same temperature setting and number of repetitions for all the LLMs.
While different temperatures may impact the output of LLMs, specific guidelines and recommendations remain general; varying from 0 to 2, where higher values indicate more random or creative generation [48].
Although more empirical studies are needed, it is common to set the temperature around 0.7:
(1) Llama 2 [39] uses 0.1 and 0.8 as defaults depending on the task and benchmark;
(2) the OpenAI API reference333https://platform.openai.com/docs/api-reference/introduction/ defaults to 0.7;
(3) open-source frameworks such as The Rasa Community444https://rasa.com/docs/rasa/next/llms/llm-intent/ use 0.7;
(4) GTP-4 uses 0.6 as their “best-guess” [32];
(5) academic studies often use 0.7, aiming to balance determinism and creativity [28, 50, 9].
Consequently, we employed 0.7 for all four models. To account for randomness, we let the LLMs generate tests for each rule times [4].
Regarding the prompt (see Section 3), only Llama2 and GPT-3.5 support separate system and user parts.
For Mistral and Mixtral, we concatenate the two parts into a single prompt.
The three models Mistral, Llama2, and Mixtral are accessed through the AI platform fireworks.ai555https://fireworks.ai/, while GPT-3.5 which is accessible only on the OpenAI platform666https://openai.com/.
4.3. Medical Rules and Rule Mutations
We selected a subset of the CRN’s rules, i.e., validation rules, of which we
excluded that contain complex structures, such as user-defined functions instead of variables, which would require access to the functions (as they are not part of the rule) and add complexity to Dvare++.
In total, we use rules in our study.
In addition, to enlarge the studied rule set in RQ1 and compute the robustness metrics in RQ2, we define a set of mutation operators on medical rules (see Table 1), inspired by previous research [36, 22].
The mutation operators consider
(1) changing logical operators (CO, NI, and ACO), set containment (RI), and built-in functions (RSE);
(2) swapping indices (SSI) and rule clauses (SR);
and
(3) altering dates (ACO).
We apply each mutation operator one at a time on every rule at every possible location.
Therefore, one mutation operator may be applied multiple times to the same rule when there are multiple possible locations, e.g., if a rule contains three and operators, we would apply CO three times.
In total, we retrieve mutated rules from the rules.
Note that we use the mutated rules only to evaluate the tests generated by the LLMs in RQ1 and RQ2 and not in RQ3, as GURI does not have these rules implemented and, hence, would not return the expected test results.
Table 1. Mutation operators
Name
ID
Description
Alter Comparison Operators
ACO
Modify > to < and <= to >= and vice versa.
Alter Date
AD
Alter the date by year, month, or day.
Change Operator
CO
Change and to or and vice versa.
Negate Inequality
NI
Change = to != and vice versa.
Reverse Inclusion
RI
Change in condition to notIn or vice versa.
Replace Startswith/Endswith
RSE
Change startswith to endswith and vice versa.
Swap Rule
SR
If Rule A implies Rule B then change it to Rule B implies Rule A.
Swap Substring Indices
SSI
If substring(i,j) then change it to substring(j,i).
4.4. Evaluation Metrics
In this section, we define the metrics used to answer the RQs.
4.4.1. RQ1
We follow previous works [7, 2, 14] and use EM to check if the completion (LLM test generation) matches exactly the expected output.
An exact match is considered when the LLM can generate syntactically correct tests for all three types, i.e., Pass, Fail, and NotApplied.
This metric allows us to assess the LLM’s ability to produce valid tests, evaluating how accurately the LLM follows the instructions in the prompt.
Considering instruction inconsistency as an indicator of potential hallucination [19], we gain insight into the model’s performance in handling different medical rules and maintaining consistency with the given instructions.
First, we introduce the completion rate metric in Eq. 1, computed as the ratio of Exact Matches (EM) to the total expected cases (), representing the overall repetitions.
We use to denote the number of times an LLM successfully generated valid tests.
Second, we study each LLM’s test generation success from two combined aspects:
(1) how successful they are in producing exactly matched outputs
and
(2) how many of the successfully matched outputs are true positives.
This is calculated as the Euclidean distance between the results of an LLM from these two aspects and the maximum possible values for them. Assuming the two points , and . represents the results for an LLM for a rule for a test type (e.g., Pass), for repetitions. is the number of times out of , the LLM produced exactly matched outputs, whereas represents how many matched outputs were true positives. Assuming , B will be . As an example—for point (25, 15) means that we observe only 25 tests out of 30 (i.e., the maximum) for a specific rule under Pass and out of 25 observations only 15 tests are actually Passing by Dvare++.
The Euclidean distance is then defined in Eq. 2.
is then normalized using min-max normalization where the minimal distance and maximal distance are calculated based on the constant coordinates for distance calculation and , respectively.
Finally, the success index for a particular rule is calculated in Eq. 3.
represents the success percentage, with higher values indicating greater success.
Derived , we denote , , and as the sets of success indices across all the rules per test type (Pass, Fail, and NotApplied).
Third, we measure the time from sending the request to the LLM platform to receiving the response, including the model inference time and other communication overheads (e.g., network).
Throughout the paper, we refer to this time as the inference time .
4.4.2. RQ2
We assess the robustness of an LLM by introducing minor changes to each rule and study how much the outputs of the original and mutated rules differ.
A similar way of assessing the robustness of LLMs has been employed by Hyun et al. [20].
To this end, we calculate the average absolute difference between the success indices of each rule in the original and mutated sets for each test type.
The result is then subtracted from 1 to represent the robustness index, where a higher value indicates greater robustness for the individual rules within the test type.
Let be the robustness index for a specific rule of test type , as defined in Eq. 4.
where
{inparablank}
4.4.3. RQ3
Regarding differential testing, we use D to denote the Dvare++ results and G to denote the GURI results. We define the metrics for a match and a mismatch as follows:
as the set of rule numbers associated with matches for a specific result, see Eq. 5.
A match is considered when both systems return the same result for the same test data.
Similarly, we define in Eq. 6 as the set of rule numbers associated with mismatches for a specific result.
4.5. Statistical Analyses
We use hypothesis testing and effect sizes to statistically evaluate our result observations.
A set of observations is a distribution of metric values, e.g., by an LLM.
Depending on the RQ, we consider different observations analyzed by the statistical tests.
We follow the best practice [4].
We use the non-parametric Kruskal-Wallis H test [25] to compare multiple observation sets, which is an extension of the pair-wise Mann-Whitney U test.
The null hypothesis states no statistical difference among the distribution sets’ ranks.
The alternative hypothesis states that there is a difference.
If we can reject and accept , we perform Dunn’s post-hoc test [13] to identify for which pairs there is a difference.
We set the significance level to and control the false discovery rate with the Benjamini-Yekutieli procedure [6] for multiple comparisons.
We additionally perform the Vargha-Delaney [41] to assess the effect size magnitude.
Two observation sets are stochastically equivalent if .
If , the first observation is stochastically better than the second observation; otherwise .
We further divide into magnitude categories, based on [18]:
negligible: ,
small: ,
medium: ,
and
large: .
All results are considered statistically significant if the p-value is below and the effect size is non-negligible.
4.6. Threats to Validity
Construct Validity
The threats to construct validity are mostly concerned with the defined metrics.
We define effectiveness in terms of completion rate and success index , and efficiency with the inference time .
Our robustness metric builds on rule mutations.
Hence, the choice of mutation operators, which is based on previous work [36, 22], is key to the validity of .
Other metrics might lead to different results.
includes the Hypertext Transfer Protocol (HTTP) request sent to the LLM platform (i.e., OpenAI or Fireworks.ai) and not only the model inference time.
However, the HTTP time should be similar across the models, as we use the same platform for Llama2, Mistral, and Mixtral, only for GPT-3.5 we use OpenAI’s API.
Internal Validity
Internal validity is about the experiment setup.
We use hosted LLMs and do not run the models ourselves, which removes our control over how the models are run.
For example, we can not be sure that model instances and sessions are not shared across users and HTTP requests, and, therefore, information from previous prompts may leak into our results.
However, we studied the platforms’ documentation and are quite certain that each request is treated as its own session and that no leakage should happen.
Another threat is the LLM temperature setting, where we picked a common value and kept it consistent across the LLMs.
The results would likely change with different temperatures.
Moreover, as LLMs are stochastic in nature, our results might also suffer from random artifacts.
To address this threat, we follow the best practice from search-based software engineering and repeat our experiment times and use appropriate statistical tests [4].
Using different statistical procedures to draw conclusions might lead to slightly different results.
Finally, every piece of software is prone to bugs; so is our experimentation pipeline.
We carefully reviewed our scripts and results and reran the experiments multiple times to ensure we do not experience critical bugs.
External Validity
Threats to external validity relate to the generalizability of our results, in particular, beyond the LLMs, medical rules, mutation operators, and the rule engine GURI.
We select a representative set of LLMs that have shown strong performance in general-purpose tasks and widespread popularity in the AI community.
In terms of rules, we select a subset of GURI’s validation rules.
Our results might change for different validation rules, aggregation rules, or different rules from other cancer registries, potentially from other countries.
Finally, we perform a laboratory experiment, i.e., we run the tests against a standalone version of GURI in an isolated setting, and do not perform a field experiment, where the tests are executed against the real GURI.
5. Results and Analyses
In this section, we discuss the results for the RQs.
5.1. RQ1: Effectiveness and Efficiency
This section investigates the completion rate , success index , and inference time across the LLMs.
Figure 3. Completion rates for all the rules per LLM.
Table 2. Statistical tests of the completion rates, success indices per test type, and inference times per LLM
Metric
Model 1
Model 2
Comp.
p-value
Magnitude
Mistral
Llama2
worse
large
Mistral
Mixtral
equal
negligible
Mistral
GPT-3.5
worse
large
Llama2
Mixtral
better
large
Llama2
GPT-3.5
worse
medium
Mixtral
GPT-3.5
worse
large
Mistral
Llama2
worse
medium
Mistral
Mixtral
worse
small
Mistral
GPT-3.5
worse
large
Llama2
Mixtral
better
medium
Llama2
GPT-3.5
worse
small
Mixtral
GPT-3.5
worse
large
Mistral
Llama2
worse
small
Mistral
Mixtral
worse
medium
Mistral
GPT-3.5
worse
large
Llama2
Mixtral
equal
negligible
Llama2
GPT-3.5
worse
small
Mixtral
GPT-3.5
worse
small
Mistral
Llama2
worse
medium
Mistral
Mixtral
worse
small
Mistral
GPT-3.5
worse
large
Llama2
Mixtral
better
small
Llama2
GPT-3.5
worse
medium
Mixtral
GPT-3.5
worse
large
Mistral
Llama2
better
large
Mistral
Mixtral
equal
negligible
Mistral
GPT-3.5
better
large
Llama2
Mixtral
worse
large
Llama2
GPT-3.5
better
small
Mixtral
GPT-3.5
better
large
Completion Rate
Figure 3 depicts the across the LLMs.
We observe that GPT-3.5 performs the best with a mean of %, indicating that it almost never hallucinates and always generates the correct JSON test output.
The second-best model is Llama2 with %, before there is a considerable drop for Mixtral and Mistral to % and %, respectively.
The statistical tests, depicted in Table 2, support these observations: GPT-3.5 is statistically better than Llama2 with medium and Mixtral and Mistral with large effect sizes.
Figure 4. Success indices for all the rules per test type and LLM.
Success Index
Figure 4 shows the success indices , , for the Pass, Fail, and NotApplied tests, respectively, across the LLMs.
Similar to the completion rates, we observe that GPT-3.5 performs the best, with a mean of %, of %, and of %,
followed by Llama2, Mixtral, and Mistral.
The statistical tests confirm GPT-3.5’s superiority: for all three test types, it is better than all the other LLMs, at least with a small effect size.
While the LLMs perform similarly on and , they exhibit a considerable drop on , except Mixtral which performs similarly across the three metrics.
This shows that the LLMs struggle more to generate Fail tests, which might be due to the structure of the rules, where the LLM must satisfy the left clause of the implication while it must falsify the right clause (see Section 2).
Figure 5. Inference times for all the rules per LLM.
Inference Time
In terms of efficiency, Fig. 5 plots the inference times for all the rules across the LLMs.
We observe that the most effective LLM, i.e., GPT-3.5, is the least efficient one, with a mean of .
The fastest LLMs are Mixtral with and Mistral with , followed by Llama2 with .
In terms of statistical tests, Mixtral and Mistral are equal, while GPT-3.5 is outperformed by Mixtral and Mistral with large and by Llama2 with small effect size.
This shows that higher effectiveness comes at a considerable runtime cost.
5.2. RQ2: Robustness
Figure 6. Robustness indices for all the rules per test type and LLM.
Table 3. Statistical tests of the robustness indices for all the rules per test type and LLM
Metric
Model 1
Model 2
Comp.
p-value
Magnitude
Mistral
Llama2
equal
negligible
Mistral
Mixtral
equal
negligible
Mistral
GPT-3.5
worse
large
Llama2
Mixtral
equal
negligible
Llama2
GPT-3.5
worse
medium
Mixtral
GPT-3.5
worse
medium
Mistral
Llama2
better
large
Mistral
Mixtral
equal
negligible
Mistral
GPT-3.5
better
medium
Llama2
Mixtral
worse
large
Llama2
GPT-3.5
equal
small
Mixtral
GPT-3.5
better
medium
Mistral
Llama2
better
large
Mistral
Mixtral
equal
small
Mistral
GPT-3.5
equal
negligible
Llama2
Mixtral
worse
large
Llama2
GPT-3.5
worse
large
Mixtral
GPT-3.5
equal
negligible
This section investigates the LLM robustness per rule and test type.
Figure 6 depicts the robustness indices , , and for the Pass, Fail, and NotApplied tests per rule, respectively.
The results are more diverse compared to the success indices: no LLM is superior across the three test types.
In terms of , GPT-3.5 is most robust with a mean of %, followed by Mixtral, Llama2, and Mistral with %, %, and %, respectively.
This is confirmed by the statistical tests (see Table 3): GPT-3.5 is better than Mixtral and Llama2 with medium and Mistral with large effect size.
For , Mistral and Mixtral tied for the most robust LLM with %, followed by GPT-3.5 with % and a medium effect size and Llama2 with % and a large effect size.
Finally for , we do not observe statistical differences between Mistral, Mixtral, and GPT-3.5 with %, %, and %, respectively.
Only Llama2 is inferior to the other three LLMs with % and large effect sizes.
5.3. RQ3: Differential Testing
We investigate matches and mismatches between Dvare++ and GURI.
In terms of mismatches, we categorize them based on the result type (i.e., Pass, Fail, NotApplied, and Warning) and “errors” returned by GURI.
A Warning result occurs when the rule execution is successful, but the data exhibits potential issues. For example, a patient’s age might be 120 years, which is theoretically possible but highly unlikely; hence the Warning result.
In addition, we identify two other result types:
(1) 500—thrown as an exception by GURI in response to invalid or non-compliant variables (e.g., date data type), deviating from the specific implementation standards. For instance, if a date is provided in a format other than expected (e.g., YYYY-MM-DD), it triggers a 500 Internal Server Error.
(2) Empty Response—similar to the above error, an empty response is triggered in specific instances where certain variables dictate conditions or requirements for the rule’s execution.
For example, if the diagnosis’ certainty, identified with the “ds” variable, is either invalid or null in a cancer message, the system generates an empty response due to pre-aggregation constraints.
We have six categories based on the GURI result types.
A mapping between GURI with Dvare++ results is done to determine if a match or mismatch is observed.
Table 4 shows all the observed matches and mismatches after executing each generated test.
We notice that Mixtral has the highest number of matches, while Mistral has the highest number of mismatches per rule. A total of 44 rule mismatches between the GURI and Dvare++ results for Mistral, followed by GPT-3.5 with 32, Llama2 with 29, and Mixtral with 27.
In addition, we perform a manual investigation of the nature of the mismatched rules and find the following patterns:
(1) Rules that contain pre-aggregation constraints.We observe that all 11 rules containing the “ds” variable result in an Empty Response when its value is invalid or missing.
(2) Rules that return 500.We encounter two rules, specifically V47 and V53, that trigger a 500 result due to either an invalid date or a non-compliant date format.
(3) Rules that return Warning.Among all the 58 rules, we observe only one rule, i.e., V08, which returns Warning instead of Fail.
(4) Rules that always return NotApplied.We observe that the rules V43, V44, and V45 always return NotApplied, regardless of the test type (i.e., Pass, Fail, NotApplied).
(5) Rules that never Pass.We observe a list of rules where older versions can never Pass. Specifically, V19/1, V20/1, V22/1, V27/1 (resulting always in Fail) and V69/1 (resulting always in NotApplied).
Based on the identified result mismatches, we observe the following reasons why such mismatches can occur:
(1) Inconsistency in handling rule versions.We notice that some of the rules have both versions “active” (e.g., V09/1 and V09/2), while some other rules are “inactive” for the first version. Thus, such rules never Pass.
(2) Inconsistency in handling date formats.We notice that GURI considers YYYY-MM-DD as a valid date format, while DD-MM-YYYY is invalid, thus leading to a 500 result.
(3) Inconsistency in handling variable dependencies.We notice that different variables are handled differently in terms of execution. For instance, the “ds” variable affects the whole cancer message (resulting in an empty response), while other variables (e.g., diagnose date) affect only specific rules (resulting always in NotApplied).
Table 4. Matches and mismatches between GURI and Dvare++ results, considering all 58 rules
Model
GURI Result
Dvare++ Result
Mismatched Rules
Match
Mismatch
Pass
44 (%)
14 (%)
V19/1, V19/2, V20/1, V20/2, V21, V25, V26, V27/1, V28, V29/1, V46, V50, V51, V67
Fail
47 (%)
11 (%)
V17, V22/1, V22/2, V23, V24, V25, V27/1, V29/1, V46, V50, V51
Mistral
Not Applied
53 (%)
5 (%)
V43, V44, V45, V46, V69/1
Warning
57 (%)
1 (%)
V08
500
57 (%)
1 (%)
V53
Empty Response
46 (%)
12 (%)
V17, V23, V24, V25, V26, V28, V47, V49, V51, V53, V67, V65
Pass
52 (%)
6 (%)
V19/2, V20/1, V20/2, V27/1, V46, V50,
Fail
52 (%)
6 (%)
V19/1, V20/1, V22/1, V22/2, V27/1, V50
Llama2
Not Applied
54 (%)
4 (%)
V43, V44, V45, V69/1,
Warning
57 (%)
1 (%)
V08
500
57 (%)
1 (%)
V53
Empty Response
47 (%)
11 (%)
V17, V23, V24, V25, V26, V28, V47, V49, V51, V65, V67
Pass
55 (%)
3 (%)
V22/2, V27/1, V46
Fail
53 (%)
5 (%)
V19/1, V22/1, V22/2, V27/1, V46
Mixtral
Not Applied
54 (%)
4 (%)
V43, V44, V45, V69/1
Warning
57 (%)
1 (%)
V08
500
56 (%)
2 (%)
V47, V53
Empty Response
46 (%)
12 (%)
V17, V23, V24, V25, V26, V28, V47, V49, V51, V53, V65, V67
Pass
52 (%)
6 (%)
V20/2, V46, V50, V51, V64, V67
Fail
49 (%)
9 (%)
V19/1, V20/1, V22/1, V22/2, V24, V26, V27/1, V46, V50
GPT-3.5
Not Applied
53 (%)
5 (%)
V43, V44, V45, V46, V69/1
Warning
57 (%)
1 (%)
V08
500
57 (%)
1 (%)
V53
Empty Response
48 (%)
10 (%)
V17, V24, V25, V26, V28, V47, V51, V53, V65, V67
6. Discussion and Lessons Learned
In this section, we discuss our learned lessons.
6.1. Hallucination Effect
As an indicator of potential hallucination, we consider the LLMs accuracy in following the given instructions in the prompt.
We utilize the metric as a binary evaluation to determine whether the tests generated by the LLMs are correct.
Despite its common use, we identify several limitations:
(1) Semantic alterations.This limitation refers to small semantic changes such as transforming “violation_case” to “violating_case”. While seemingly minor, such a change is considered incorrect.
(2) Different levels of detail.We observe discrepancies in the level of detail between the expected and generated tests. This manifests in three common scenarios:
(a) Missing test types: In some instances, the LLMs fail to generate all specified test types. For example, out of the three expected test types (Pass, Fail, and NotApplied), only two might be generated (Pass and Fail).
(b) Additional tests: The LLMs might produce redundant tests, unnecessarily generating variations of the same scenario (e.g., generating “invalid_case”, “invalid_case_1”, and “invalid_case_2”).
(c) Lack of integration: Tests are generated individually and not as a single JSON dictionary containing all three types.
(3) Invalid JSON format.The JSON parser throws exceptions for issues including:
(a) Improperly formatted property names: JSON requires property names to be enclosed in double quotes.
(b) Missing variable and value pairs: Errors can occur when the JSON misses information, such as variable names or their corresponding values.
(c) Incorrect data structure: Instead of generating the expected dictionary format (key-value pairs), the LLM generates lists.
(d) Missing delimiters: JSON parsing errors also occur when expected delimiters, e.g., “:” or “,”, are missing.
6.2. No LLM Subsumes Another
Our differential testing results reveals that the most effective LLM (i.e., GPT-3.5) does not uncover the most mismatches.
Surprisingly, it is Mistral that detects the highest number of mismatches.
However, it is important to note that not all mismatches indicate faults.
We identify three specific cases:
(1) inconsistency in handling rule versions,
(2) inconsistency in handling date formats,and
(3) inconsistency in handling variable dependencies.
While 22 rules fall into these categories, none of the LLMs manages to identify all of them.
For instance, while Mistral successfully detects all rules involving the “ds” variable, it overlooks cases triggering a 500 status code—unlike Mixtral.
Contrarily, Mixtral identifies both cases triggering a 500 status code but misses those where all the rule versions (e.g., V20/1) never Pass.
Interestingly, both Llama2 and GPT-3.5 detect all the cases related to rule versions but fail to uncover the cases where Mistral and Mixtral excel.
While any of the LLMs would be suited for detecting mismatches, the optimal choice for a developer or researcher requires deeper understanding of the specific inconsistency types each model excels at identifying and the concrete usage scenario.
Additionally, it is important to remain open to exploring other LLMs that potentially improve on detecting specific inconsistencies or offer novel insights.
6.3. Using LLMs Instead of Other Test Generation Techniques
A specialized testing approach is essential for ensuring the correctness of the medical rules and their implementation in GURI.
While common testing methods like random and coverage-guided testing are widely used [17, 51], they fall short of capturing the specifics of medical rules, such as rule context and relationships among variables.
LLMeDiff introduces a customized, domain-aware strategy that leverages LLMs to generate tests tackling the unique aspects of medical rules and GURI (e.g., ICD codes).
Moreover, as the medical rules are similar to constraints, a test generation technique might rely on constraint solvers to find suitable tests.
While such solvers are efficient in finding satisfying and violating variable assignments, they may not capture the “meaning” of variable names and relationships among variables and constraints. Nonetheless, a dedicated empirical evaluation is needed to compare LLMs with other constraint-solving approaches (e.g., based on search algorithms and satisfiability modulo theories (SMT) solvers).
6.4. LLMs Might Not Find Simple Faults
The power of LLMs is that they can interpret the variable names to generate valid values.
For dates this can be helpful, as a date string needs to follow a specific format (e.g., YYYY-MM-DD); otherwise, the rule engine will yield a parsing error.
An LLM will always generate a date string of a valid format; however, it might never try a simple input such as a single number (e.g., 7).
This specific case would be detected in our differential testing setting, where Dvare++ yields an error but GURI accepts a single number as a date variable, but the LLM never generates such a simple case.
A simpler test generator, e.g., a fuzzer like EvoMaster [3], would likely try such an input and detect the fault.
6.5. Mismatches Faults
In RQ3, we discussed the differential testing results in terms of rule result matches and mismatches.
Our study revealed that LLMeDiff serves as an effective technique for uncovering disparities within the rule engines.
However, it is important to emphasize that LLMeDiff’s task is limited to flagging these mismatches.
There is a potential for false positives and negatives (e.g., data overlaps between rule versions—erroneously leading to Pass), which complicates the interpretation of the results.
False positives might stem from the reference implementation (Dvare++) not accurately reflecting the desired behavior of GURI.
Therefore, determining whether these differences signify real faults in the rule engine requires manual root cause analysis by the CaReSS developers.
While such false positives add work for the developers, we assume that checking whether a mismatch is a false positive will neither require much time nor effort, as a domain expert can quickly identify a false positive.
Over time, the reference implementation can be adapted to handle edge cases that previously led to false positives.
Alternatively, it is worth investigating whether learned models can serve as a reference implementation, e.g., a digital twin of GURI [29].
6.6. Improvement on the State of the Practice
The rule testing process at the CRN is currently done manually by medical coders, who run a set of tests against new or changed rules whenever “necessary” [27].
Instead, automated rule test generation would relieve them of this tedious task.
Unfortunately, and to the best of our knowledge, there are no automatic test generation techniques for medical rule engines that have a solution for the oracle problem [5].
In our previous works, we explicitly pointed out that generating valid medical data [26], solving the oracle problem for medical rules and rule engines [26], and keeping the number of tests executed against the real system to a minimum [21] is detrimental for being applicable at the CRN.
In this paper, we address these challenges by generating only three valid and realistic tests, i.e., Pass, Fail, and NotApplied, for each rule and using differential testing with a simple reference implementation, i.e., Dvare++.
6.7. Generality of the Results
While the results are specific to the CRN, CaReSS, and GURI; the medical rules used in the experiment; and the studied LLMs; we believe that LLMeDiff and the overall findings are transferable to other contexts.
(1) Our findings are relevant to other health registries in Norway (e.g., rthe Norwegian Patient Registry) and similar real-world contexts, such as the Norwegian toll and customs, which implement similar rule-based systems.
(2) Other countries also maintain health registries, such as cancer registries, for which our findings and lessons can provide valuable guidance for adopting and applying an LLM-based test generation approach such as LLMeDiff.
(3) Our medical rules are an instance of logic-based rules or constraints; therefore, any system taking such rules as input can leverage an approach like LLMeDiff.
(4) The differential testing approach of LLMeDiff uses a simple reference implementation to compare against, as alternative implementations of highly customized systems, such as GURI, are unlikely to exist.
Hence, systems with similar settings aiming to test beyond simple crash oracles can follow a testing approach like ours.
7. Related Work
A typical software testing challenge in many contexts is the absence of precise and reliable test oracles [8], i.e., checking whether a test case’s outcome meets the expected outcome using a test oracle. In many real-world situations (including our context of testing medical rule engines), such an oracle either doesn’t exist or is impractical due to resource constraints. Metamorphic and differential testing techniques have been employed in the literature to deal with the test oracle problem [35, 15]. In this paper, we applied a typical differential testing setup to support testing GURI, i.e., using Dvare++ as a simplified reference implementation of GURI. The key idea is to use GURI and Dvare++ to check whether both systems produce consistent outputs when provided identical inputs generated by LLMs. Such a comparison allows for identifying inconsistencies, hence identifying possible issues in the implementation of GURI. As a result, we address the test oracle problem to some extent.
Recently, LLMs have become popular in software testing [43], e.g., focusing on unit testing [16, 47], system input generation [12, 11], property-based test generation [42], and test oracle generation [40]. Our work is similar to these works, focusing on generating tests with LLMs that let medical rules yield Pass, Fail, or NotApplied results; however, the application context is specific to the medical domain.
Consequently, our work is related to the works that aim to explore the potential of LLMs in the medical domain. For example,
Zhang et al. [49] evaluates the performance of LLMs, including GPT-3.5, GPT-4, Falcon, and LLama 2, in the classification of medical terms, specifically using a metastatic cancer dataset to assess reasoning consistency. Wornow et al. [46] evaluates models such as GPT-4 on clinical text and concludes that many models are trained on small or broad biomedical datasets, emphasizing the importance of a better evaluation system that aligns with healthcare standards. In general, surveys [34, 38, 10] provide insights into the potential and challenges of LLMs in healthcare applications and highlight that the practical use of LLMs remains limited. Our work with the CRN is a step towards the practical deployment of LLMs for testing their software systems.
In the past, we developed a model-based framework to model the medical rules for GURI [44] with the Unified Modeling Language (UML) and Object Constraint Language (OCL). Next, we developed an impact analysis approach [45] to study the impact of rule changes in response to GURI’s evolution (e.g., changes in regulations). In addition, we developed a search-based approach to automatically refactor rules [30] to enhance rule understandability and maintainability. However, these approaches do not help finding problems in the modeled rules. More recently, four studies advanced the testing at the CRN, mainly focusing on the GURI subsystem [26, 27, 21, 29].
Laaber et al. [27] address the practical complexities of testing an evolving system, offering insights applicable to healthcare registries. Moreover, a case study focused on applying EvoMaster [3] (an AI-based testing tool) to assess its effectiveness regarding code coverage, error detection, and domain-specific rule coverage at the CRN [26].
Meanwhile, we also proposed EvoClass [21] by extending EvoMaster with a ML classifier, showing significant cost reductions without compromising effectiveness. Lastly, EvoCLINICAL [29], a cyber-cyber digital twin for GURI, was proposed, which with fine-tuning through active transfer learning, can deal with system evolution. In contrast to these existing works, we make new contributions by addressing the test oracle problem associated with rule execution results by employing differential testing. Moreover, we tackle the challenge of generating electronic health records (EHRs) by leveraging the capabilities of LLMs.
8. Conclusions
This paper presented LLMeDiff, an LLM-based test generation approach for medical rules employed to perform differential testing of GURI, the CRN’s medical rule web service.
While we find that LLMeDiff is effective in generating tests and finding mismatches, there are a few future directions worth investigating:
(1) generate functions with LLMs that produce Pass, Fail, and NotApplied tests;
(2) using LLMs to generate rule mutations;
(3) fine-tune an LLM with domain-specific EHRs and CRN data;
and
(4) combine other test generation techniques with LLMeDiff.
Data Availability
Due to confidentiality, we can not disclose the data, scripts, and systems that lead to the results of the paper.
Acknowledgements.
The research presented in this paper has received funding from The Research Council of Norway (RCN) under the project 309642 and has benefited from the Experimental Infrastructure for Exploration of Exascale Computing (eX3), which is supported by the RCN project 270053.
The Norwegian Ministry of Education and Research supports Erblin Isaku’s PhD.